OpenStudy (anonymous):

Need help solving, I am horrible at word problems :) In a recent year, total sales at Rest A was $400,000 more than Rest B. Total sales combined is $550,000,000. What were the sales for each store? Rest A=? Rest B=?

jimthompson5910 (jim_thompson5910):

Start by defining variables

jimthompson5910 (jim_thompson5910):

Let x = total sales at Rest A and y = total sales at Rest B

jimthompson5910 (jim_thompson5910):

Since "total sales at Rest A was $400,000 more than Rest B", we can say x = y + 400000

OpenStudy (anonymous):

ok, im following

jimthompson5910 (jim_thompson5910):

and because "Total sales combined is $550,000,000", we also know that x+y = 550,000,000

jimthompson5910 (jim_thompson5910):

So start with x+y = 550,000,000 and plug in x = y+400,000 x+y = 550,000,000 (y+400,000)+y = 550,000,000 y+400,000+y = 550,000,000

jimthompson5910 (jim_thompson5910):

Now solve for y

OpenStudy (anonymous):

ok, give me a sec to solve and check my answer please :)

jimthompson5910 (jim_thompson5910):

ok, will do

OpenStudy (anonymous):

y=274,800,000?

jimthompson5910 (jim_thompson5910):

that is correct

jimthompson5910 (jim_thompson5910):

now use this value of y to find the value of x

OpenStudy (anonymous):

x+2748000000=550000000?

jimthompson5910 (jim_thompson5910):

yes, now solve for x

OpenStudy (anonymous):

x=219,800,000 Rest A= $219,800,000 Rest B= $274,800,000

OpenStudy (anonymous):

but Rest A should be more

jimthompson5910 (jim_thompson5910):

Rest A is off a bit

OpenStudy (anonymous):

should I subtract 400,000 from rest B and add to rest A?

jimthompson5910 (jim_thompson5910):

oh i see what went wrong

jimthompson5910 (jim_thompson5910):

x+2748000000=550000000 should be x+274800000=550000000 you put one extra zero in 274800000

jimthompson5910 (jim_thompson5910):

x+274800000=550000000 x=550000000-274800000 x=275,200,000

OpenStudy (anonymous):

I did? My mistake. Sorry

jimthompson5910 (jim_thompson5910):

no worries

jimthompson5910 (jim_thompson5910):

there are too many zeros to keep track of anyway

OpenStudy (anonymous):

thats the truth...lol

jimthompson5910 (jim_thompson5910):

it might have been better to let x = total sales of rest A (in thousands of dollars) and y = total sales of rest B (in thousands of dollars) but it's too late for that

OpenStudy (anonymous):

no worries. I got what I needed and thanks for your help and breaking it down for mw

jimthompson5910 (jim_thompson5910):

np